If one year of college currently costs $20,000, how much will one year cost in 18 years for a newborn son assuming an inflation rate of 4%?
Answer #1


solution: FV=PV*(1+i)^n FV=20000*(1.04118) FV PV i Future value Present value inflation rate number of years FV= 40,516.33